OpenStudy (jenniferjuice):

Solve for x: |x − 2| + 10 = 12

OpenStudy (jenniferjuice):

x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No Solution

OpenStudy (johnweldon1993):

Well with absolute values...you first want to get it alone so you have |x - 2| + 10 = 12 you want the absolute value alone...so subtract 10 from both sides... this will leave you with |x - 2| = 2 Now this gives you 2 different equations.... x - 2 = 2 and x - 2 = -2 solve both of those equations...and you will have your answer(s)

OpenStudy (jenniferjuice):

um is it .... x=-1 and x= 1 ?

OpenStudy (jenniferjuice):

or maybe x=4 and x= 0 ?

OpenStudy (johnweldon1993):

Not quite.... our 2 equations are x - 2 = 2 and x - 2 = -2 so lets look at the first one x - 2 = 2 Remember we want to solve for 'x'....since 2 is being taken away from 'x'....ask yourself....WHAT minus 2....equals 2? And for the other one x - 2 = -2 same process....WHAT minus 2....equals -2?

OpenStudy (jenniferjuice):

wait would i add ?

OpenStudy (jenniferjuice):

x - 2 = 2 + 2 +2 x = 4

OpenStudy (jenniferjuice):

@Loser66 hey sorry to bother but can you help me?

OpenStudy (loser66):

the first one, because if you have x=0, |0-2| =2 then +10 =12 check x =4 ,|4-2|=2 then +10 =12 check therefore, the first one is the right answer

OpenStudy (jenniferjuice):

so x = 0 and x = 4 ?

OpenStudy (loser66):

yes

OpenStudy (johnweldon1993):

Yeah sorry about getting off....but yeah good job!